OpenStudy (anonymous):

i need algebra help! okay (12q-5)(m+3) how do we do it and why is the answer 12mq-5m+36q-15???

OpenStudy (amistre64):

if you know how to multiply 2 number together (10)(5) it is the same concept here. First off; Do you know what the distributive property is?

OpenStudy (anonymous):

nope my mother doesnt know what it is im helpin my mother actually and ik what it is but its hard to explain what it is for me

OpenStudy (amistre64):

the distributive propery simply states that we can multiply number together. Lets take an easy example: (3)(8) = 24 , this we should already know. 3(5+3) should equal the same thing right? because (5+3) = 8 lets distribute (3) into (5+3) like this: 3(5) + 3(3) = 15 + 9 = 24 Does that make sense?

OpenStudy (anonymous):

yeah,whats the explaining of the steps?

OpenStudy (amistre64):

If you want to know why you get what you got for an answer, the steps need to be explained, right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

we can take this property one step further like this: (1+2)(5+3) should equal 24..its the same numbers. what we get is: (1+2)(5) + (1+2)(3) now we distribute again to get: 5(1) + 5(2) + 3(1) + 3(2) = 5+10+3+6 = 24 Does that make sense? cause if it doesnt, then we cant figure out your original problem :)

OpenStudy (anonymous):

no not really :(

OpenStudy (amistre64):

Your original problem is: (12q-5)(m+3) lets distribute like this: (12q-5)(m) + (12q-5)(3) now distibute one more time to get: 12q(m) -5(m) + 12q(3) -5(3) 12qm -5m +36q -15 is what we end up with but you have to understand what is going on to get to it. There are shortcuts taht you can use, but they dont teach you WHY it works.

OpenStudy (amistre64):

let me know what doesnt make sense, and maybe I can be of more help :)

OpenStudy (anonymous):

okay thankx u've been a lot help